# Learning Prompt-Enhanced Context Features for Weakly-Supervised Video Anomaly Detection

### v1

@@ Abstract
Video anomaly detection under weak supervision is challenging due to the absence of frame-level annotations during the training phase. Previous work has employed graph convolution networks or self-attention mechanisms to model temporal relations, along with multiple instance learning (MIL)-based classification loss to learn discriminative features. However, most of them utilize multi-branches to capture local and global dependencies separately, leading to increased parameters and computational cost. Furthermore, the binarized constraint of the MIL-based loss only ensures coarse-grained interclass separability, ignoring fine-grained discriminability within anomalous classes. In this paper, we propose a weakly supervised anomaly detection framework that emphasizes efficient context modeling and enhanced semantic discriminability. To this end, we first construct a temporal context aggregation (TCA) module that captures complete contextual information by reusing similarity matrix and adaptive fusion. Additionally, we propose a prompt-enhanced learning (PEL) module that incorporates semantic priors into the model by utilizing knowledge-based prompts, aiming at enhancing the discriminative capacity of context features while ensuring separability between anomaly sub-classes. Furthermore, we introduce a score smoothing (SS) module in the testing phase to suppress individual bias and reduce false alarms. Extensive experiments demonstrate the effectiveness of various components of our method, which achieves competitive performance with fewer parameters and computational effort on three challenging benchmarks: the UCF-crime, XD-violence, and ShanghaiTech datasets. The detection accuracy of some anomaly sub-classes is also improved with a great margin.
@@ I Introduction
Video anomaly detection (VAD) refers to the identification of uncommon events or behaviors that deviate from typical patterns in video streams [1, 2, 3]. Given the growing prevalence of surveillance cameras, human supervision alone is insufficient to fulfill the demands of practical applications. As such, the need for efficient and accurate automated VAD methods has rapidly increased in recent years. These methods have become increasingly important in various domains, including security and public safety, industrial process monitoring, healthcare, and social media analysis. By preventing potential threats and improving operational efficiency, VAD methods hold significant practical value.
In contrast to classical action recognition tasks, collecting relevant videos for anomaly detection poses a challenge due to the sensitivity and scarcity of anomalous events. The mainstream approach treats anomaly detection as a semi-supervised learning task [4], wherein the characterization patterns of normal data are learned during the training phase, while identifying samples that deviate from the normality as anomalies [5, 6, 7, 8, 9, 10]. While this paradigm does not necessitate fine-grained annotation, the non-exhaustive nature of normal patterns often results in the identification of unseen samples as anomalies, leading to a high false alarm rate.
In recent years, there has been a significant focus on developing more robust algorithms for anomaly detection, with particular attention given to weakly supervised approaches [11, 12, 13, 14, 15]. Compared to semi-supervised methods, weakly supervised paradigm offers several advantages: (1) The training data includes both normal and abnormal videos, enabling the model to learn more discriminative representations. (2) Weak supervision only requires video-level annotations, where the presence of anomalous events is indicated, but their start and end locations are unknown. (3) The absence of fine-grained annotations allows for the construction of large-scale datasets for anomaly detection. Thanks to these advantages, the weakly supervised approach has achieved superior performance and substantially outperforms the semi-supervised approach.
Considering that the duration of anomalous events can vary significantly, it is crucial to model the temporal relationship of video snippets to capture contextual information. Currently, graph convolutional networks [16] and self-attention mechanisms[17] are widely used for temporal relation modeling. However, the former only aggregates nodes within a finite neighborhood of video snippets, ignoring the integrity of anomalous events, while the latter primarily captures global correlations of all clips, which may introduce redundant noise. To address these limitations, Wu et al. [18] proposed an HL-Net with two paralleled graph convolutional networks to capture holistic and local dependencies, respectively, and obtained fused features through a simple concatenation operation. Similarly, Tian et al. [13] constructed an MTN network that combines multi-head self-attention and pyramid dilated convolutions to capture multi-scale temporal relations. These methods effectively consider local-global context modeling, thereby suppressing the interference of long-range noise. However, such parallel structure of multiple branches inevitably increases the parameters and computational cost, which may result in over-fitting.
Figure 1: Complex abnormal events from different acquisition sources. (a) UCF-Crime [11]. (b) XD-Violence [18].
In addition, the diversity and complexity of anomalies result in large intra-class variance and small inter-class variation within anomaly datasets, as depicted in Figure 1. Such datasets commonly include anomalous events in surveillance videos or in film and TV works. The former is captured by fixed-view cameras while the latter contains rich artistic expressions such as camera changes and scene switching. Building a robust anomaly detection pipeline in these complex scenarios is a challenging task. To address this issue, most studies construct a compact normal manifold with additional constraints. Wu et al. [12] introduced a center loss and a margin ranking loss to learn a discriminative decision boundary. Similarly, Park et al. [19] used learnable vectors to construct a set of prototypes for normal samples and introduced a normality-guided clustering loss with a triplet loss to enhance the inter-class separability between abnormal and normal snippets. However, these methods overlook the discriminability within anomaly classes, as illustrated in Figure 1, where there are significant differences between arson and fighting, burglary and shooting at the visual level, even though they are all anomalous events. Moreover, such binarized constraints fail to learn the specific semantics of anomalies, resulting in poor interpretability of predictions. Therefore, we contend that learning fine-grained context features for video anomaly detection is a critical aspect that facilitates the model to make more reliable decisions.
In this paper, we aim to address the two key issues mentioned above, i.e., temporal context modeling and limited anomaly discriminability, to further improve the anomaly detector. Firstly, we propose a temporal context aggregation (TCA) module to capture temporal relations across video snippets. Unlike existing parallel architectures in [13] [18] [20] [21], the TCA module enables efficient context modeling by reusing the similarity matrix. We also use a learnable factor for adaptive fusion of local-global contexts and embed a dynamic position encoding (DPE) into the similarity matrix, which uses a parametric kernel function to model the relative location of video snippets. Our TCA module outperforms existing methods with fewer parameters and computational cost, as demonstrated by experimental results in Section IV.
Secondly, inspired by prompt learning [22], we propose a prompt-enhanced learning (PEL) module to strengthen the fine-grained semantics of context features, as shown in Figure 2. Existing methods [23, 24, 25, 26] use hand-crafted or learnable vectors to construct prompt templates. The former are mostly class-agnostic phrases lacking sufficient context information, while the latter are abstract strings with weak interpretability. In contrast, we construct prompt templates by extracting relevant concepts from an external knowledge base, i.e., ConceptNet [27], which can serve as effective supplements to the context while possessing a certain degree of interpretability. By aligning anomalous contexts with their corresponding prompt features and repulsing non-anomalous contexts, rich semantics are injected into visual features to enhance their fine-grained discriminability. This operation also enlarges the inter-class separability between positive and negative snippets, which is conducive to forming a more discriminative decision boundary in the embedding space.
Finally, to address possible misclassifications caused by factors such as frame jitter or switching, we introduce a score smoothing (SS) module in the testing phase to suppress individual bias by applying average pooling to consecutive anomaly scores. This effectively improves the performance of our model while reducing false alarms in normal videos.
The main contributions of this paper are summarized as follows:
@@
    We construct a TCA module by reusing the similarity matrix and adaptive fusion for video anomaly detection, which achieves superior encoding performance with fewer parameters and computational cost.
@@
    We propose a PEL module that enables fine-grained context enhancement by constructing knowledge-based prompts, fore-background separation and cross-modal alignment.
@@
    Our model achieves competitive results on three benchmarks, i.e., UCF-Crime, XD-Violence, and ShanghaiTech dataset. We also observe a significant improvement in the detection accuracy of some fine-grained anomalies of approximately 10%.
The rest of the paper is organized as follows: Section II provides a review of the existing work on video anomaly detection and prompt learning. Section III presents the proposed method and its details. Section IV showcases the comparative results of state-of-the-art methods on three benchmarks and validates the effectiveness of each component through ablation studies. Finally, Section V concludes the paper.
Figure 2: Comparison of different methods for prompt construction. (a) Hand-crafted prompt. (b) Learnable prompt. (c) PEL module with knowledge-based prompt (Ours). The context features are enhanced by cross-modal alignment with the corresponding prompt features.
@@ II-A Video Anomaly Detection
As previously noted, early work [28, 29, 30, 31] considered anomaly detection as a semi-supervised learning task, i.e., only normal samples are available during the training phase, where statistical models with hand-crafted features were utilized. However, these approaches were limited by the weak representation ability of such features, leading to poor robustness and generalization. In recent years, the emergence of deep models has enabled the development of more powerful anomaly detection algorithms. Some approaches [32, 33, 34, 35] aim to construct a one-class classifier that learns a compact manifold by constraining the latent space of normality, while identifying samples that deviate from that manifold as anomalies. However, due to the absence of a priori knowledge of anomalous events, some normal samples that are complex or unseen may be erroneously classified as anomalies. Other methods [36, 37, 38] assume that normality can be well reconstructed or predicted and identify samples with larger reconstruction or prediction errors as anomalies. Video anomaly detection commonly employs deep models such as U-Net [39], generative adversarial networks (GAN) [40], and ConvLSTM [41]. However, these methods are prone to over-fitting the training data, resulting in the reconstruction or prediction of some anomalies as normal events.
Recently, several studies have focused on weakly supervised anomaly detection, which involves training with both normal and abnormal samples that have video-level annotations. Sultani et al. [11] proposed a deep multi-instance learning model that utilizes a margin ranking loss to maximize the anomaly scores between positive and negative snippets. However, since the ranking loss requires paired inputs, Wan et al. [42] developed a dynamic multi-instance loss employing a top-k strategy. Subsequently, Tian et al. [13] introduced the Robust Temporal Feature Magnitude (RTFM) learning method to detect slight anomalies based on feature similarity and temporal continuity. The RTFM method parallelizes the self-attention block and pyramid dilated convolution layers to capture multi-scale temporal dependencies. Li et al. [43] proposed the Multi-Sequence Learning (MSL) method to refine localization boundaries by progressively optimizing a reduced sampled length. The self-attention mechanism incorporates Depth-wise Convolution to lower computational complexity for sequence correlation. Other studies, such as GCN-Cleaner [20] and MIST [44], generate high-confidence pseudo-labels for video snippets, thereby transforming weakly supervised anomaly detection into a supervised learning task under noisy labels.
@@
Although some of these approaches have investigated temporal relation modeling, many rely on parallel branches that require additional parameters and computations. In contrast, our method offers two distinct advantages. Firstly, the TCA module achieves local-global context encoding by reusing the similarity matrix, thereby outperforming methods that require extra branches in [13] [18] [21] [45] while using fewer parameters and computations. Secondly, our PEL module injects anomaly priors into the model through external knowledge instead of constructing the normality prototype from within the data as in [12] and [19]. This approach enhances the semantics of visual features and further improves the fine-grained anomaly detection performance of the model.
@@ II-B Prompt Learning in Video Understanding
Prompt Learning was originally proposed in the field of natural language processing with the goal of leveraging pre-trained models to achieve favorable generalization in few-shot or zero-shot scenarios via the design of suitable prompt templates. Recent research has successfully applied this approach to video understanding tasks. Wang et al. [24] aligned video clips with corresponding prompt embeddings of category labels for action recognition, incorporating category labels as prefix, fill-in-the-blank, and suffix into prompt templates to achieve superior performance over the single category prompt. Ju et al. [26] constructed prompt templates by combining category labels with a set of learnable vectors and explored the impact of label position in the templates. However, designing prompts manually is a time-consuming task and the model is sensitive to the content of the template [46]. While parametric prompts do not require linguistic priors, the learned prompts are often abstract strings with limited interpretability.
Yao et al. [47] utilized the definitions provided by WordNet [48] to create prompt templates. They also generated complementary descriptions of category labels by constructing a concept dictionary, which yielded outstanding results in open-vocabulary object detection tasks. Inspired by their work, we propose a novel prompt-enhanced learning approach for video anomaly detection. Our method differs from theirs in that we extract class-specific concepts highly relevant to anomalies from ConceptNet [27] to enrich potential context information. Furthermore, we aim to enhance the visual features by incorporating external prompts. To achieve this, we inject semantic information into visual features through cross-modal alignment, rather than bidirectional matching, which ensures that only visual features are required during the testing phase.
@@ III-A Overall Framework
The overall framework of our method is shown in Figure 3. Specifically, an untrimmed video is first divided into non-overlapping snippets by a sliding window of 16 frames. Subsequently, we use a pre-trained I3D network [49] to extract the snippet features $\mathbf{X}$, where the subscripts $a$ and $n$ denote abnormal and normal videos, respectively, which are fed into the TCA module to generate context features $\mathbf{X}^{c}$. After that, feature reduction is achieved through a two-layer multilayer perceptron (MLP), with the PEL module applied to the middle layer to learn discriminative features $\mathbf{X}^{s}$ via knowledge-based prompts. Finally, the CC module predicts snippet-level anomaly scores $\mathbf{S}$. During the training phase, the MIL-based loss function converts the snippet-level scores into bag-level predictions to learn high activations for anomalous cases. In the testing phase, the SS module is utilized to suppress individual biases.
Figure 3: Overview of our proposed framework. The model comprises six components in addition to the feature extraction network (I3D). The TCA module captures both local and global context simultaneously and facilitates adaptive fusion. The high-level representations are obtained using a two-layer MLP, with the PEL module operating in the middle layer to learn fine-grained semantics for context features. The CC module functions as a classifier to predict snippet-level anomaly scores. In the training phase, the model is optimized using both $\mathscr{L}_{ce}$ and $\mathscr{L}_{kd}$, while SS is utilized for bias suppression during inference.

### v2

@@ Abstract
Video anomaly detection under weak supervision presents significant challenges, particularly due to the lack of frame-level annotations during training. While prior research has utilized graph convolution networks and self-attention mechanisms alongside multiple instance learning (MIL)-based classification loss to model temporal relations and learn discriminative features, these methods often employ multi-branch architectures to capture local and global dependencies separately, resulting in increased parameters and computational costs. Moreover, the coarse-grained interclass separability provided by the binary constraint of MIL-based loss neglects the fine-grained discriminability within anomalous classes. In response, this paper introduces a weakly supervised anomaly detection framework that focuses on efficient context modeling and enhanced semantic discriminability. We present a Temporal Context Aggregation (TCA) module that captures comprehensive contextual information by reusing the similarity matrix and implementing adaptive fusion. Additionally, we propose a Prompt-Enhanced Learning (PEL) module that integrates semantic priors using knowledge-based prompts to boost the discriminative capacity of context features while ensuring separability between anomaly sub-classes. Extensive experiments validate the effectiveness of our method’s components, demonstrating competitive performance with reduced parameters and computational effort on three challenging benchmarks: UCF-Crime, XD-Violence, and ShanghaiTech datasets. Notably, our approach significantly improves the detection accuracy of certain anomaly sub-classes, underscoring its practical value and efficacy. Our code is available at: https://github.com/yujiangpu20/PEL4VAD.
@@ I Introduction
Video Anomaly Detection (VAD) is the process of identifying unusual events or behaviors that diverge from normal patterns in video streams [1, 2, 3]. With the increasing ubiquity of surveillance cameras, relying solely on human supervision is inadequate for meeting the demands of practical applications. Consequently, there’s a burgeoning need for efficient and accurate automated VAD methods. These methods are crucial across various sectors, including security, industrial monitoring, healthcare, and social media analysis, where they help mitigate potential threats and enhance operational efficiency, thereby holding significant practical value.
Unlike classical action recognition tasks, gathering relevant videos for anomaly detection is challenging due to the rarity and sensitive nature of anomalous events. The prevailing approach treats anomaly detection as a semi-supervised learning task [4], where the model learns the patterns of normal behavior during training and identifies deviations as anomalies [5, 6, 7, 8, 9, 10]. These approaches don’t require detailed annotations but often mistakenly identifies unseen samples as anomalies due to the incomplete representation of normal patterns, resulting in high false alarm rates.
Recently, there’s been a notable shift toward developing more robust algorithms for anomaly detection, with a specific focus on weakly supervised methods [11, 12, 13, 14, 15]. These offer several advantages over semi-supervised approaches: 1) They incorporate both normal and abnormal videos in training, enabling more discriminative representation learning. 2) They require only video-level annotations, indicating the presence but not the exact timing of anomalies. 3) They facilitate the creation of large-scale datasets due to less stringent annotation requirements. Owing to these benefits, weakly supervised methods have demonstrated superior performance, significantly outshining their semi-supervised counterparts.
To effectively model the temporal dynamics of anomalous events, which vary significantly in duration, it’s crucial to understand the contextual information within video snippets. Existing methods, such as graph convolutional networks [16] and self-attention mechanisms [17], aim to capture these temporal relationships. However, each has its limitations: graph convolutions only consider local node aggregates, potentially missing the broader context of anomalies, while self-attention mechanisms, despite capturing global correlations, might introduce irrelevant noise. To refine this understanding, Wu et al. [18] and Tian et al. [13] devised networks that integrate local and global contexts through parallel structures. Although these approaches enhance context modeling by mitigating long-range noise interference, they come at the cost of increased computational complexity and parameter count, which can lead to overfitting and hinder practical deployment.
Figure 1: Comparison of different methods for prompt construction. (a) Hand-crafted prompt. (b) Learnable prompt. (c) Knowledge-based prompt (Ours). The context features are enhanced by cross-modal alignment with the corresponding prompt features.
Furthermore, the inherent diversity and complexity of anomalies pose additional challenges. Anomaly datasets often showcase large intra-class variance and minimal inter-class variation, with examples ranging from fixed-camera surveillance footage to artistically varied film and TV content. Addressing this requires a robust detection pipeline capable of navigating these complexities. While recent methods aim to construct a compact normal manifold [12] and introduce additional binary constraints [19] for discriminative learning, they often neglect the distinct visual characteristics within anomaly classes and fail to capture the specific semantics of various anomalies. This oversight leads to poor interpretability and an increased risk of misclassification.
In light of these limitations, our motivation is twofold: 1) to develop an approach that balances the need for detailed temporal modeling with computational efficiency, thereby addressing the high parameter count and low efficiency issues inherent in multi-branch structures. 2) to construct a method that recognizes and interprets the diversity and complexity of anomaly semantics more effectively than existing models.
In this paper, we firstly introduce a Temporal Context Aggregation (TCA) module designed to capture temporal relations across video snippets more efficiently than the current parallel architectures [13] [18] [20] [21]. This module distinguishes itself by reutilizing the similarity matrix, thereby reducing the computational load and parameter count. It also incorporates a learnable factor to adaptively fuse local and global contexts and introduces a Dynamic Position Encoding (DPE) with a parametric kernel function to accurately model the relative positions of video snippets. As evidenced by the results in Section IV, our TCA module demonstrates superior performance with a more streamlined and cost-effective design compared to existing methods.
Secondly, we propose a Prompt-Enhanced Learning (PEL) module, drawing inspiration from prompt learning [22]. Unlike existing methods [23, 24, 25, 26] that rely on hand-crafted or learnable prompts, our approach leverages an external knowledge base, i.e., ConceptNet [27], to construct prompt templates, as demonstrated in Figure 1. These templates are contextually rich and possess a degree of interpretability, offering a nuanced understanding of the specific semantics of anomalies. By aligning anomalous contexts with corresponding prompt features and distancing non-anomalous contexts, our PEL module injects rich semantics into visual features, enhancing fine-grained discriminability and inter-class separability. This innovative approach fosters the development of a more discerning decision boundary in the embedding space.
Finally, considering that anomalies in videos generally extend across several frames, we utilize a plug-and-play score smoothing (SS) strategy during the testing phase to address potential mis-classifications due to transient disturbances such as frame jitter or switching. This module applies average pooling to consecutive anomaly scores, thereby suppressing individual biases and mitigating the impact of ephemeral noise inconsistent with the typical duration of an anomaly.
The main contributions of this paper are succinctly outlined as follows:
@@
    We introduce a Temporal Context Aggregation (TCA) module that leverages a reused similarity matrix and adaptive fusion for efficient and effective video anomaly detection. This module demonstrates enhanced encoding performance while significantly reducing the parameter count and computational expense.
@@
    We present a Prompt-Enhanced Learning (PEL) module designed to augment fine-grained contextual understanding. This is achieved through the construction of knowledge-based prompts, precise context separation, and cross-modal alignment, thereby enriching the semantic discrimination of the model.
@@
    Our model exhibits competitive performance across three benchmark datasets: UCF-Crime, XD-Violence, and ShanghaiTech. Notably, it shows an approximate 10% improvement in detection accuracy for certain fine-grained anomalies, underscoring its practical efficacy.
The remainder of the paper is structured as follows: Section II reviews related work in video anomaly detection and prompt learning. Section III delineates the proposed method in detail. Section IV presents comparative results with state-of-the-art methods on benchmark datasets and validates the effectiveness of each component through ablation studies. Finally, Section V concludes the paper, reflecting on findings and future directions.
@@ II-A Video Anomaly Detection
Early studies [28, 29, 30, 31] treated anomaly detection as a semi-supervised task, utilizing statistical models with hand-crafted features. These approaches, however, suffered from limited representational capacity, leading to suboptimal robustness and generalization. Recent advancements in deep learning have spurred the development of more sophisticated anomaly detection techniques. Some methods [32, 33, 34, 35] aim to create a one-class classifier that delineates normality in a latent space, identifying deviations as anomalies. Yet, without prior knowledge of anomalies, these models may mislabel complex or unseen normal samples as anomalous. Other strategies [36, 37, 38, 39] postulate that anomalies manifest as larger reconstruction or prediction errors, assuming normality can be accurately reconstructed or predicted. However, these methods are prone to overfitting and may reconstruct or predict anomalies well.
Recent research has pivoted towards weakly supervised anomaly detection, leveraging both normal and abnormal samples with video-level annotations. These methods generally categorize into two paradigms: one-stage methods utilizing Multiple Instance Learning (MIL) and two-stage self-training strategies. Sultani et al. [11] introduced a deep MIL model optimizing a margin ranking loss to enhance the differentiation between positive and negative snippets. Tian et al. [13] developed Robust Temporal Feature Magnitude (RTFM) learning, detecting subtle anomalies through feature similarity and temporal continuity. Cho et al. [40] devised Class-Activate Feature Learning (CLAV) to extract distinctive features and employed relative distance learning to accentuate their differences, alongside the Context-Motion Interrelation Module (CoMo) for analyzing scene-related motion anomalies. Conversely, Lv et al. [41] proposed an unbiased MIL approach that considers both confident and ambiguous snippets, enhancing the detector’s ability to differentiate between normal and abnormal events without relying on contextual biases.
Two-stage self-training methods have emerged to generate high-confidence pseudo-labels for video snippets, recasting weakly supervised anomaly detection as a supervised task with noisy labels. Zhong et al. [20] employed a graph convolution network to refine these labels and iteratively enhance the anomaly classifier. Feng et al. [42] proposed utilizing multi-instance pseudo label generation to fine-tune feature encoders, thereby yielding task-specific discriminative features. Li et al. [43] introduced Multi-Sequence Learning (MSL), a method that incrementally optimizes reduced sample lengths to refine localization boundaries. Zhang et al. [15] introduced a multi-head module to generate varied pseudo labels and an iterative uncertainty-based training strategy that refines the process by prioritizing clips with lower uncertainty.
In parallel, several studies have leveraged multimodal information to enhance video anomaly detection. Wu et al. [18] proposed HL-Net, integrating appearance, motion, and audio for a comprehensive multimodal approach. Yu et al. [44] utilized a dual-stream network to improve instance clustering and employ self-distillation to minimize the semantic gap between unimodal and audio-visual features. Wei et al. [45] introduced the MSAF framework, employing multimodal MIL ranking loss to generate pseudo clip-level labels and a supervise-attention regression loss in feature learning to foster implicit alignment between different modalities. Pu et al. [46] enhanced contextual video representation through cross-modal interactions between audio and visual elements, utilizing temporal convolution to compute high-confidence scores for specific events like violence.
@@
While numerous methods have explored temporal relation modeling, many depend on parallel branches that introduce additional parameters and computational demands. In contrast, our approach presents two significant advantages. Firstly, the TCA module capitalizes on the reuse of the similarity matrix to concurrently capture local-global dependencies. This strategy outperforms methods requiring extra branches [13], [18], [21], and [47], while minimizing both parameters and computational load. Secondly, diverging from techniques that construct normality prototypes from the data itself, like [12] and [19], our PEL module introduces anomaly priors into the model using external knowledge. This approach not only enriches the semantics of visual features but also notably advances the fine-grained anomaly detection capabilities of our model.
@@ II-B Prompt Learning in Video Understanding
Prompt learning, initially introduced in natural language processing, aims to utilize pre-trained models for few-shot or zero-shot scenarios by designing appropriate prompt templates. This technique has recently been extended to video understanding tasks. Wang et al. [24] demonstrated this by aligning video clips with prompt embeddings of category labels for action recognition. They incorporated category labels as prefixes, fill-in-the-blanks, and suffixes into prompt templates, thereby significantly outperforming single category prompts. Similarly, Ju et al. [26] crafted prompt templates by merging category labels with learnable vectors, examining the impact of label positioning within the templates. Despite these advancements, manual prompt design remains laborious, and models are highly sensitive to template content [48]. While parametric prompts eliminate the need for linguistic priors, they often result in abstract and cryptic strings.
Recently, Yao et al. [49] leveraged WordNet’s [50] definitions to formulate prompt templates and created additional category label descriptions using a concept dictionary. This method significantly advanced open-vocabulary object detection tasks. Inspired by their work, we introduce a novel prompt-enhanced learning strategy for video anomaly detection. Unlike previous methods, we extract anomaly-relevant, class-specific concepts from ConceptNet [27] to augment contextual information. Our objective is to improve visual features through external prompts. We achieve this by infusing semantic information into visual features via cross-modal alignment, rather than bidirectional matching, ensuring reliance solely on visual features during testing.
@@ III-A Overall Framework
The overall framework of our method is shown in Figure 2. Specifically, an untrimmed video is first divided into non-overlapping snippets by a sliding window of 16 frames. Subsequently, we use a pre-trained I3D network [51] to extract the snippet features $\mathbf{X}$, which are fed into the TCA module to generate context features $\mathbf{X}^{c}$. After that, feature reduction is achieved through a two-layer multilayer perceptron (MLP), with the PEL module applied to the middle layer to learn discriminative features $\mathbf{X}^{e}$ via knowledge-based prompts. Finally, the classifier predicts snippet-level anomaly scores $\mathbf{S}$. During the training phase, the MIL-based loss function converts the snippet-level scores into bag-level predictions to learn high activations for anomalous cases, where the subscripts $a$ and $n$ denote abnormal and normal videos, respectively.
Figure 2: Overview of the proposed framework. We first process untrimmed videos using a pre-trained I3D network to extract snippet features. The TCA module then simultaneously captures both local and global contexts. High-level representations are derived using a two-layer MLP, with the PEL module operating in the middle layer to learn fine-grained semantics for context features. Finally, a causal convolution layer functions as a classifier to predict snippet-level anomaly scores. During the training stage, the model is optimized using both cross entropy loss $\mathscr{L}_{ce}$ and KL-divergence loss $\mathscr{L}_{kd}$.